OpenStudy (thecatman):

Which verbal phrase represents the expression 1 - one minus nine-tenths n n? "one minus nine-tenths of a number" "the difference one minus nine-tenths" "nine-tenths of a number minus one" "nine-tenths the difference of one minus a number"

OpenStudy (thecatman):

sorry ignore the first n

OpenStudy (anonymous):

\[1-(1-\frac{ 9 }{ 10 }n)\]

OpenStudy (anonymous):

is n n mean \[1-(1-\frac{ 9 }{ 10 }n) n \] ?

OpenStudy (anonymous):

the first one

OpenStudy (thecatman):

\[1-\frac{ 9 }{ 10 } n\]

OpenStudy (anonymous):

yes